# Symmetric functions and the explicit moment problem for abelian groups

### v1

@@ 3. Proofs and extensions
#$\#\operatorname{Sur}(G_{\lambda},G_{\mu})=\frac{P_{\lambda/\mu}(t,t^{2},\ldots;0,t)}{P_{\lambda}(t,t^{2},\ldots;0,t)Q_{\mu}(1,t,\ldots;0,t)}.$ (3.1)
@@
when $\mu^{\prime}\succ\nu^{\prime}$, and is 00 otherwise. Combining (3.4) with (3.5) shows (1.3). Uniqueness is clear from (1.3). ∎
Given a random abelian group $G$, it is often desirable to study the quotient $G/p^{d}G$. For instance, in [Woo17] and subsequent works on cokernels of random integer matrices, one always reduces the cokernel modulo some integer $a$, for then one only has to study random matrices over a finite ring $\mathbb{Z}/a\mathbb{Z}$. 1.1 admits the following refinement, which concerns the $p^{d}$-torsion group $G/p^{d}G$ and consequently allows $G$ to be potentially infinite (though its $p$-part must be finitely generated for the moment finiteness condition to be satisfied). In what follows we let $\mathbb{Y}_{d}^{\prime}=\{\lambda\in\mathbb{Y}:\operatorname{len}(\lambda^{\prime})\leq d\}$.
@@
One may also show 3.2 by noting that for $p^{d}$-torsion $H$, #$\#\operatorname{Sur}(G,H)$ depends only on $G/p^{d}G$, and then doing similar manipulations to (3.4) with a sum over $\mu_{d}^{\prime}$. The dependence of the moments on $G/p^{d}G$ may be seen directly in symmetric function theory (though this would be a very convoluted way to arrive at it): By the explicit formulas 2.3 and their generalization [VP22a, Theorem 3.3] to skew Hall-Littlewood functions, for any $\lambda\in\mathbb{Y},\mu\in\mathbb{Y}_{d}^{\prime}$ we have
@@
which is independent of $\lambda_{d+1}^{\prime},\lambda_{d+2}^{\prime},\ldots$. In the special case of 3.2 given in [VP23, Lemma 4.2], we were not yet aware of the simpler proof of 3.2 via 1.1 and basic group theory given above, and so argued directly in this way similarly to (3.4) using (3.7).
@@
For any finite collection $P=\{p_{1},\ldots,p_{k}\}$ of primes, we let $\mathcal{A}_{P}$ the set of (isomorphism classes of) finite abelian groups $H$ such that every prime factor of #$\#H$ lies in $P$. We further define, for any partitions $\lambda(1),\ldots,\lambda(k)\in\mathbb{Y}$, the group
$G_{\lambda(1),\ldots,\lambda(k)}(P):=\bigoplus_{i=1}^{k}G_{\lambda^{(i)},p_{i}}.$ (3.8)
It is clear that $\mathcal{A}_{P}=\{G_{\lambda(1),\ldots,\lambda^{(|P|)}}(P):(\lambda(1),\ldots,\lambda^{(|P|)})\in\mathbb{Y}^{|P|}\}$.
@@
It is equivalent to show the case of constant $G$, i.e. the equality
@@
of functions on $\mathcal{A}_{P}$, as the general case may be obtained by linear combinations of the above. We do this by induction on $k$, the base case $k=1$ following from 1.1. Suppose (3.10) holds for some $k-1$ and let $p_{1},\ldots,p_{k}$ be distinct primes. Writing
@@
#$\#\operatorname{Sur}(G,G_{\mu(1),\ldots,\mu(k)}(P))=\#\operatorname{Sur}\left(\bigoplus_{i=1}^{k-1}G_{p_{i}},\bigoplus_{i=1}^{k-1}G_{\mu(i),p_{i}}\right)\cdot\#\operatorname{Sur}(G_{p_{k}},G_{\mu(k),p_{k}}).$ (3.12)
@@
The same proof as for 3.2 yields a version of the above formula for $\Pr(G/(p_{1}^{e_{1}}\cdots p_{k}^{e_{k}})\cong H)$ in the setting of 3.3. It is this version for modules over the finite quotient ring $\mathbb{Z}/(p_{1}^{e_{1}}\cdots p_{k}^{e_{k}})$ which corresponds to the formula in [SW22b, Proposition 6.2], though one must do some calculation with that formula to reduce it to ours.
@@
If $R$ is a complete discrete valuation ring with maximal ideal $(\omega)$ generated by a uniformizer $\omega$, and finite residue field $R/(\omega)\cong\mathbb{F}_{q}$, then 1.1 and 3.2 hold with $\operatorname{Sur}(G,H)$ replaced by surjective $R$-module homomorphisms and $t=p^{-1}$ replaced by $t=q^{-1}$. One only needs that 3.1 holds in this setting, which follows because the number of surjections is known to be a rational function of $q^{-1}$ in this setting, see for instance [Mac98, Chapter II].

### v2

@@ 3. Proofs and extensions
$\operatorname{Sur}(G_{\lambda},G_{\mu})=\frac{P_{\lambda/\mu}(t,t^{2},\ldots;0,t)}{P_{\lambda}(t,t^{2},\ldots;0,t)Q_{\mu}(1,t,\ldots;0,t)}.$ (3.1)
@@
when $\mu^{\prime}\succ\nu^{\prime}$, and is 0 otherwise. Combining (3.4) with (3.5) shows (1.3). Uniqueness is clear from (1.3). ∎
Given a random abelian group $G$, it is often desirable to study the quotient $G/p^{d}G$. For instance, in [27] and subsequent works on cokernels of random integer matrices, one always reduces the cokernel modulo some integer $a$, for then one only has to study random matrices over a finite ring $\mathbb{Z}/a\mathbb{Z}$. 1.1 admits the following refinement, which concerns the $p^{d}$-torsion group $G/p^{d}G$ and consequently allows $G$ to be potentially infinite (though its $p$-part must be finitely generated for the moment finiteness condition to be satisfied). In what follows we let $\mathbb{Y}_{d}^{\prime}=\{\lambda\in\mathbb{Y}:\operatorname{len}(\lambda^{\prime})\leq d\}$.
@@
One may also show 3.2 by noting that for $p^{d}$-torsion $H$, $\operatorname{Sur}(G,H)$ depends only on $G/p^{d}G$, and then doing similar manipulations to (3.4) with a sum over $\mu_{d}^{\prime}$. The dependence of the moments on $G/p^{d}G$ may be seen directly in symmetric function theory (though this would be a very convoluted way to arrive at it): By the explicit formulas 2.3 and their generalization [24, Theorem 3.3] to skew Hall-Littlewood functions, for any $\lambda\in\mathbb{Y},\mu\in\mathbb{Y}_{d}^{\prime}$ we have
@@
which is independent of $\lambda_{d+1}^{\prime},\lambda_{d+2}^{\prime},\ldots$. In the special case of 3.2 given in [26, Lemma 4.2], we were not yet aware of the simpler proof of 3.2 via 1.1 and basic group theory given above, and so argued directly in this way similarly to (3.4) using (3.7).
@@
For any finite collection $P=\{p_{1},\ldots,p_{k}\}$ of primes, we let $\mathcal{A}_{P}$ the set of (isomorphism classes of) finite abelian groups $H$ such that every prime factor of #H\#H lies in $P$. We further define, for any partitions $\lambda(1),\ldots,\lambda(k)\in\mathbb{Y}$, the group
$G_{\lambda(1),\ldots,\lambda(k)}(P):=\bigoplus_{i=1}^{k}G_{\lambda(i),p_{i}}.$ (3.8)
It is clear that $\mathcal{A}_{P}=\{G_{\lambda(1),\ldots,\lambda(|P|)}(P):(\lambda(1),\ldots,\lambda(|P|))\in\mathbb{Y}^{|P|}\}$.
@@
It suffices to show the case of constant $G$, i.e. the equality
@@
of functions on $\mathcal{A}_{P}$, as the general case may be obtained by linear combinations of the above. We do this by induction on $k$, the base case $k=1$ following from 1.1. Suppose (3.10) holds for some k1k-1 and let $p_{1},\ldots,p_{k}$ be distinct primes. Writing
@@
$\operatorname{Sur}(G,G_{\mu(1),\ldots,\mu(k)}(P))=\#\operatorname{Sur}\left(\bigoplus_{i=1}^{k-1}G_{p_{i}},\bigoplus_{i=1}^{k-1}G_{\mu(i),p_{i}}\right)\cdot\#\operatorname{Sur}(G_{p_{k}},G_{\mu(k),p_{k}}).$ (3.12)
@@
The same proof as for 3.2 yields a version of the above formula for $\Pr(G/(p_{1}^{e_{1}}\cdots p_{k}^{e_{k}})\cong H)$ in the setting of $3.$ It is this version for modules over the finite quotient ring $\mathbb{Z}/(p_{1}^{e_{1}}\cdots p_{k}^{e_{k}})$ which corresponds to the formula in [22, Proposition 6.2], though one must do some calculation with that formula to reduce it to ours.
@@
If $R$ is a complete discrete valuation ring with maximal ideal $(\omega)$ generated by a uniformizer $\omega$, and finite residue field $R/(\omega)\cong\mathbb{F}_{q}$, then 1.1 and 3.2 hold with $\operatorname{Sur}(G,H)$ replaced by surjective $R$-module homomorphisms and $t=p^{-1}$ replaced by $t=q^{-1}$. One only needs that 3.1 holds in this setting, which follows because the number of surjections is known to be a rational function of $q^{-1}$ in this setting, see for instance [14, Chapter II].